a 164-inch board is cut into two pieces. One piece is three times the length of the other. Find the lengths of the two pieces?
Answer by jim_thompson5910(35256)

first piece: x

second piece: 3x (3 times longer than the first)

They add to 164 inches

x + 3x = 164

4x = 164

x = 164/4

x = 41

3x = 3*41 = 123

Therefore, the two pieces are 41 inches and 123 inches